GO003A

Session: May 2000 (session 2)
Wavelength: 18 cm

Comment


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Zacchiroli at Sat May 27 23:11:02 METDST 2000) Success

Torun (K.Borkowski at Sat May 27 23:19:10 METDST 2000) Success
No problems, although not smoothly wound tapes may result in a poorer playback. Weather predominantly fine (day sunny) till tonight evening when an atmospheric storm passed nearby.

Noto (C.Trigilio at Mon May 29 10:26:53 METDST 2000) Success

Effelsberg (A. Kraus at Mon May 29 11:29:55 METDST 2000) Minor failures
At the beginning a wrong tape label was given. Since this was not rejected by the system, during the first five scans a wrong tape speed and footage was set. Only part of the data was recorded. Scans 3 and 5 were lost. This was corrected at UT 2235 (day 147). The failure of the field system computer caused the loss of data from UT 1912-2015 and UT2035-2100 (day 148).

Westerbork (Tony Foley at Tue May 30 10:31:30 METDST 2000) Minor failures
Used full array (14 telescopes). From 21:00 to 21:48 polarization was swapped. Lost scans from 03:08 to 06:46 with DCB computer not initializing, and again from 12:47 to 14:03. Had to reboot systems. Several scans from 14:05 to 16:29 out of hour angle range, so not observed.

Westerbork (Tony Foley at Tue May 30 10:34:55 METDST 2000) Minor failures
A hacker was active in WSRT during this period. There was no obvious interference in the online systems.

Jodrell_Bank (Alastair Gunn at Mon Jun 5 12:23:26 METDST 2000) Minor failures
Usual RFI at 1672 and 1685 MHz. Receiver performance looks very good except for ringing around these RFI spikes. Running with IF filter centred on 1658 MHz and 16 MHz bandwidth to keep interference out. Problems with telescope control may have caused some scans to be lost.